The safety air bags in automobiles are inflated by nitrogen gas generated by the rapid decomposition of sodium azide, NaN3 (molar mass= 65.01 g/mol). If an air bag has a volume of 36 L and is to be filled with nitrogen gas at a pressure of 1.15 atm at a temperature of 26.0 °C, how many grams of NaN3 must be decomposed? _____ g NaN3


1
Expert's answer
2022-05-24T15:49:03-0400

The reaction of azide decomposition:

"2NaN_3\u21922Na+3N_2"


The mole number can be deduced from the given parameters:

"pV=nRT" ,

where p is a pressure in atm, V is the volume in litres (L), n is a mole number, R is a gas constant ("0.082 \frac{L\times atm}{R \times mole}" ) and T is the temperature in Kelvin.


"n=\frac{pV}{RT} = \frac{1.15 atm \times 36 L}{0.082 \times (26+273)K} = 1.69 mole"


The mass of azide can be obtained from mole number:

"m = nM = 1.69 mole \times 65.01 g/mol = 109.78 g"